the altitude of an isosceles trapezoid is 20cm. if the lower base is 10cm longer than the upper base the area of the trapezoid is 220sq. cm, find the lengths of the bases.: the altitude of an isosceles trapezoid is 20cm. if the lower base is 10cm longer than the upper base the area of the trapezoid is 220sq. cm, find the lengths of the bases.
Answer by jojo14344(282)

Area of a trapezoid is similar to a triangle, the only difference it has 2 bases. If a triangle is (1/2)(h)(b), we need to add another base to calculate for the area of trapezoid -----> (1/2)h(a+b)
.
In the problem, let "a"=upper base="x", "b"=lower base="x+10cm", and
altitude=h=20cm
Substituting in our eqn with given, A=220 sqcm},
220 =(1/2)(20)(x+x+10) = (1/2)(20)(2x+10)
cross multiply,
440=40x+200
440-200=40x
240=40x
x=6cm = upper base
10+6=16cm = lower base
In doubt? to go back to our eqn and you'll see, A=(1/2)(h)(a+b)
thank you,
Jojo